a no. has two digits whose sum is 14. if 36 is added to the no., its digit are interchanged . find the no.
Answer by ptaylor(1894)

Let x=the tens digit
And let y=the units digit
The number then would be (10x+y)
Now we are told that:
x+y=14------------------------eq1
and
(10x+y)+36=10y+x subtract 10x and also y from each side
10x+y-10x-y+36=10y-y+x-10x collect like terms
36=9y-9x or
-9x+9y=36------------------eq2
multiply eq1 by 9(we get 9x+9y=126) and then add it to eq2:
18y=162 divide each side by 18
y=9-------------------------------the units digit
substitute y=9 into eq1:
x+9=14 subtract 9 from each side
x=14-9
x=5--------------------the tens digit
CK
5+9=14
14=14
and
59+36=95
95=95

Hope this helps---ptaylor